Discrete Math

Use division into two cases to prove that for every integer n, 2\n(3n + 1).
Answer #1

hemce, 2/2K(64+1) Solution : To prove that evey inleger in n(3n+1) as we are asked to division into Now, two cases, Case 1 inCase 2:7 n = odd Inlegen. - 2n-1 (NEN) Now 3M (301 n3n+1] (2n-1) (31 24-1) + 3) = (21-1) (64-3 +1) = (2n-1)(64-2) = QM-1) 213